# Bidding in Ancillary Service Markets: An Analytical Approach Using Extreme Value Theory

### v2

@@ IV Distribution fitting and validation
This section outlines the methodology for fitting a probability distribution to our data, providing the necessary input for the analytical solution presented in Section V.
IV-A Fitting a Weibull distribution to the tail data
@@
Figure 3: Distribution of the downwards flexibility at the arbitrarily selected hour 13. The blue distribution is constructed using in-sample data, while the pink part represents the out-of-sample data. A Weibull distribution with the shape parameter $\hat{\gamma}=0.9265$ is fitted to the lower tail of the blue distribution. The vertical dotted line indicates the 10th percentile value of the in-sample data, below which the Weibull distribution is fitted.
For each flexibility $\{R^{\uparrow},R^{\downarrow},R^{\rm{E_{20}}}\}$, we select a subset of realizations for each hour for which we fit a distribution to the tail. As illustrated in Fig. 2, the relevant tail, in line with the case study and the P90 requirement, consists of all data points below the 10th percentile, $r\leq r_{\varepsilon}=r_{0.1}$, where we let $r$ denote a realization of an arbitrary flexibility. Since the data in this tail does not resemble a Weibull distribution, we transform the data using $x=-r+r_{0.1}$, to which the Weibull distribution is more appropriately fitted. Fig. 3 illustrates the distribution of realizations of the estimated downwards flexibility, $r^{\downarrow}$, at hour 13. In line with Fig. 2, the blue bars represent the distribution based on in-sample data, while the pink bars correspond to the remaining out-of-sample data. We aim to fit a distribution based on the in-sample data. The threshold of the in-sample distribution, below which the Weibull distribution is fitted, i.e., the 10th percentile of the sampled realizations, is indicated by a vertical dotted line.
We begin the procedure of fitting the Weibull distribution [17] by defining its two-parameter probability density function (PDF) as follows:
$f(x)$ =$=\begin{cases}\kappa\gamma x^{\gamma-1}\exp(-\kappa x^{\gamma}),&x\geq 0,\\ 0,&x<0,\end{cases}$ (3a)
@@
$\ell_{n}(x;\hat{\kappa}(\hat{\gamma}),\hat{\gamma})$ $=n\left(\log n-\log\left(\sum_{i}x_{i}^{\hat{\gamma}}\right)+\log\hat{\gamma}-1\right)+$
@@
which we maximize over a range of $\gamma$-values. Looking back at Fig. 3, the Weibull distribution is fitted to the lower tail of the sampled data, with the shape parameter $\hat{\gamma}=0.9265$. As it is needed in later analysis, we define the cumulative distribution function (CDF) of the Weibull distribution as:
$F(x)$ =$=\begin{cases}1-\exp(-\kappa x^{\gamma}),&x\geq 0,\\ 0,&x<0.\end{cases}$ (4a)
@@
IV-B Validation of distribution fits
To validate whether the fitted distribution correctly reflects the data, we use the Kolmogorov-Smirnov (KS) test [18]. This is a nonparametric test used to compare the equality of continuous, one-dimensional probability distributions, and it tests whether a sample comes from a specified reference distribution. The test statistic reports the maximum absolute difference between the empirical CDF, $F_{n}(x)$, and the reference CDF, $F(x)$, as:
@@ IV-B Validation of distribution fits
where the empirical CDF is defined as
@@
As before, $n$ represents the number of samples included in the estimation. Using (5), we can assess whether to reject or accept the hypothesis that the realizations come from the reference distribution through hypothesis testing. In this context, the null hypothesis $H_{0}$ assumes that the data comes from the specified reference distribution, while the alternative hypothesis $H_{1}$ assumes that it does not.
@@
To further compare the goodness-of-fit between several distributions, such as the fit of the Weibull distribution for a range of parameter values, the negative log-likelihood (NLL) value is utilized, where a lower value signifies a better fit. This is because the NLL effectively quantifies how well a given model and its parameters explain the observed data, with lower NLL values corresponding to higher likelihoods of the observed data under the model. In essence, we use the negative of (3d) and compare the resulting value from a range of $\gamma$-values, and choose the $\gamma$ which gives us the lowest [19]. By using the NLL and the KS test, we validate that the obtained distribution fits are satisfying for our analysis.
V Analytical solution
This section presents the proposed analytical solution for addressing the chance-constrained bidding problem introduced in Section III.
V-A Bonferroni correction
$x$$y$$\mathbb{P}(X>x,Y>y)$$\mathbb{P}(Y\leq y)$$\mathbb{P}(X\leq x)$$\mathbb{P}(X\leq x,Y\leq y)$
The joint chance constraint in (2) can be decomposed into three individual constraints using the Bonferroni correction [15]. This adjustment redistributes the overall allowance for violation equally across each constraint, effectively lowering the significance level for each test to ensure that the total probability of violation remains within the desired limit:
@@ V-A Bonferroni correction
The Bonferroni approach essentially replaces the solution space for each decision variable in (2b) with a smaller set, as presented in (7), which may make the solution more conservative. On the other hand, one might replace (2b) with a larger solution space by setting the right-hand side of the inequalities in (7) to $\varepsilon$. These sets may act as lower and upper bounds on the solution of (2), respectively. We elaborate with a two-dimensional example: using the joint probability $\mathbb{P}(x\leq X,y\leq Y)\geq 1-\varepsilon$, shaded in gray in Fig. 4, we can restate it as:
@@
Any solution of $x$ and $y$ which satisfies (8d) will then satisfy (8c) $\Rightarrow$ (8b) $\Rightarrow$ (8a). The inequality (8c) implies that threshold values of failure probability are distributed over any combination of $x$ and $y$, meaning that $x$ and $y$ may take any values satisfying $\mathbb{P}(X\leq x)<\epsilon$ or $\mathbb{P}(Y\leq y)<\epsilon$ as long as the sum of the two probabilities also satisfies (8c). The Bonferroni corrected version (8d) expresses that at most $\varepsilon/2$ of the failure can be associated with each variable, meaning that the possible values of $x$ and $y$ have a tighter constraint. If the distribution of $X$ and $Y$ are the same, while the intersect is zero, then this is the exact solution where (8d) = (8c). In Fig. 4, this is the case if we assume that the probability mass in the pink and purple shaded areas (excluding the joint probability area) are equal, i.e., $\mathbb{P}(X\leq x)=\mathbb{P}(Y\leq y)$ and $\mathbb{P}\left(X<x,Y<y\right)=0$.
V-B Analytical reformulation of (2)
Following the procedure in Section IV, the Weibull distribution is fitted to data under a threshold level, $r_{0.1}$, which in the present case is the empirical 10th percentile of each flexibility, i.e., $\mathbb{P}(R\leq r_{0.1})=\varepsilon$. Applying Bayes’ theorem, we can thus express the conditional probability for an arbitrary random variable $R$ being under a certain threshold $b$ as:
@@ V-B Analytical reformulation of (2)
[…]
VI Sample-based solution for benchmarking
@@
VI-A Sample-based reformulation of the optimization problem
@@
where $\{M^{\uparrow},M^{\downarrow},M^{\rm{E_{20}}}\}$ are sufficiently large positive constants, and $y_{i}$ is an auxiliary binary variable corresponding to sample $i$. Constraints (13b)-(13d) ensure that if $y_{i}=0$, all constraints are fulfilled for a specific sample $i$, while if $y_{i}=1$, these constraints are relaxed. Constraint (13e) ensures that the total budget for constraint violations is preserved, allowing for at most a 10% violation of the samples, in accordance with the P90 requirement. Due to the reduced number of data points in the case study, the problem remains tractable and can be solved without the need for additional relaxation.
VI-B Sampling
To find the minimum number of randomly selected samples necessary to accurately represent the estimated data, the following formula is used [20]:
@@ VI-B Sampling
VI-C Ex-post out-of-sample validation of constraints
@@
VII Numerical results
Figure 6: Illustrated in the leftmost plots are the bids for the upward and downward flexibility, respectively, using the two discussed methods: sample-based and analytical. The middle and right plots show the out-of-sample constraint violation rate, presented separately for each constraint and as an overall violation rate. All plots show the mean results obtained over 10 runs, with each run using different in-sample and out-of-sample sets, although the number of samples in each set per run remains unchanged.
In this section, we first present numerical results regarding the bidding strategy of the EV aggregator, obtained from both the analytical and sample-based solutions. We also report the rate of out-of-sample constraint violations and provide a sensitivity analysis with respect to $\varepsilon$. All source codes are publicly available in [21].
VII-A Bidding strategy
A total of 10 simulations were run, where, following (14), $|\Omega|$ samples were randomly selected at each run and used for the procedures described in Sections V and VI. As a result, average reserve capacity bids over all runs into the FCR-D up and down markets are illustrated in the top and bottom left plots of Fig. 6. In the remaining plots, the average constraint violation rate is reported based on the out-of-sample validation procedure outlined in Section VI-C, using 40% of the samples not employed in the optimization procedure.
@@ VII-A Bidding strategy
where $\pi_{h}^{(\cdot)}$ represents the prices of the FCR-D capacity reservations, and $b^{(\cdot)*}_{h}$ is the bid made at hour $h$ in either direction. Using prices from the early auction of the FCR-D up and down markets from the relevant period, meaning the period spanning that of the collected data, and the results of the reserve bids, we obtain a reservation payment of €136,292 for the sample-based method and €105,366 for the analytical method when summing over the entire year. It is of relevance to extend the formula to include penalties for bid violations, to get a full picture of the profit of the aggregator, but that is outside the scope of this paper. However, reflecting on the results regarding constraint violation, there might be clues that the bids made by using the sample-based method may result in more penalties.
VII-B Sensitivity analysis
For the sensitivity analysis, we focus solely on our proposed analytical approach.
It is of interest to study how the optimal reserve capacity bids change with varying values of $\varepsilon$, particularly how they might change when $\varepsilon$ becomes smaller, imposing a stricter requirement on reliability of the bids.
In (10), the allowance for failure is distributed across all requirements involved in the chance constraints. We can utilize this formulation to create a constraint that allows for a failure allowance lower than the level used in the distribution fitting scheme. Recall that we are using data from the lower 10th percentile to fit distributions to each flexibility. This allows us to investigate whether we can find solutions to the bidding strategy with a violation allowance lower than 10%. To perform the sensitivity analysis on the failure allowance, we can define the constraints similarly to (12b)-(12d), replacing the right-hand side of (10) with an arbitrary $\alpha\in(0.0005,0.1]$, where $\alpha=0.1/3$ corresponds to the Bonferroni corrected problem. Following the same procedure as in Section V, we obtain the optimization problem for the bidding strategy:

### v4

@@ 4 Distribution fitting and validation
This section outlines the methodology for fitting a probability distribution to our data, providing the necessary input for the analytical solution presented in Section 5.
4.1 Fitting a Weibull distribution to the tail data
@@
For each flexibility $\{R^{\uparrow},R^{\downarrow},R^{\rm{E_{20}}}\}$, we select a subset of realizations for each hour for which we fit a distribution to the tail. As illustrated in Figure 2, the relevant tail, in line with the case study and the P90 requirement, consists of all in-sample data points below the 10th percentile, $r\leq r_{\varepsilon}=r_{0.1}$, where we let $r$ denote a realization of an arbitrary flexibility. Since the data in this tail does not resemble a Weibull distribution, we transform the data using $x=-r+r_{0.1}$, to which the Weibull distribution is more appropriately fitted. Figure 3 shows the distribution of realizations of the estimated downward flexibility, $r^{\downarrow}$, for a randomly selected hour. In line with Figure 2, the blue bars represent the distribution based on in-sample data, while the pink bars correspond to the remaining out-of-sample data. We aim to fit a distribution based on the in-sample data. The threshold of the in-sample distribution, below which the Weibull distribution is fitted, i.e., the 10th percentile of the sampled realizations, is indicated by a vertical dashed line.
Figure 3: Distribution of the downwards flexibility at the arbitrarily selected hour 13. The blue distribution is constructed using in-sample data, while the pink part represents the out-of-sample data. A Weibull distribution with the shape parameter $\hat{\gamma}=0.9265$ is fitted to the lower tail of the blue distribution. The vertical dashed line indicates the empirical 10th percentile value of the in-sample data, below which the Weibull distribution is fitted.
We begin the procedure of fitting the Weibull distribution (MLE) by defining its two-parameter probability density function (PDF) as follows:
$f(x)$ $=\begin{cases}\kappa\gamma x^{\gamma-1}\exp(-\kappa x^{\gamma}),&x\geq 0,\\ 0,&x<0,\end{cases}$ (3a)
@@
[…]
@@
which we maximize over a range of appropriate $\gamma$-values. Looking back at Figure 3, the Weibull distribution is fitted to the lower tail of the sampled data, with the shape parameter $\hat{\gamma}=0.9265$. As it is needed in later analysis, we define the cumulative distribution function (CDF) of the Weibull distribution as:
$F(x)$ $=\begin{cases}1-\exp(-\kappa x^{\gamma}),&x\geq 0,\\ 0,&x<0.\end{cases}$ (4a)
@@
4.2 Validation of distribution fits
To validate whether the fitted distribution correctly reflects the data, we use the Kolmogorov-Smirnov (KS) test (KSTest). This is a nonparametric test used to compare the equality of continuous, one-dimensional probability distributions, and it tests whether a sample comes from a specified reference distribution. The test statistic reports the maximum absolute difference between the empirical CDF, $F_{n}(x)$, and the reference CDF, $F(x)$, as:
@@ 4.2 Validation of distribution fits
where the empirical CDF is defined as:
@@
As before, $n$ denotes the number of samples used in the estimation. Using (5), we can determine whether to reject or accept the hypothesis that the realizations originate from the reference distribution. In this context, the null hypothesis $H_{0}$ assumes that the data follow the specified reference distribution, whereas the alternative hypothesis $H_{1}$ assumes that they do not.
@@
To further compare the goodness-of-fit between several distributions, such as the fit of the Weibull distribution for a range of parameter values, the negative log-likelihood (NLL) value is utilized, where a lower value signifies a better fit. This is because the NLL effectively quantifies how well a given model and its parameters explain the observed data, with lower NLL values corresponding to higher likelihoods of the observed data under the model. In essence, we use the negative of (3d) and compare the resulting value from a range of $\gamma$-values, and choose the $\gamma$ which gives us the lowest (MLNLL). By using the NLL and the KS test, we validate that the obtained distribution fits are satisfying for our analysis.
4.3 Validation results
Figure 4: Best (left) and worst (right) distribution fits (grey) compared to all data below the 10th percentile (pink). The $p$-values for the best fit range from 0.871 to 0.999, while for the worst fit they range from 0.218 to 0.871.
Figure 4 presents the results of the distribution fits for the best and worst average fits over 10 runs, based on the validation metrics described in Section 4.2. Recall that each run uses a different set of in-sample data, while keeping the number of samples constant. Additionally, average results across all hours for all flexibilities are provided in Table 1 in Appendix B. The $p$-values associated with the maximum absolute difference ($D_{n}$) indicate that fits from all runs are accepted ($p>0.05$), albeit with varying strengths. For the best fit, $p$-values range from 0.871 to 0.999, with an average maximum absolute difference of 0.105, whereas for the worst fit, $p$-values range from 0.218 to 0.871, with an average maximum absolute difference of 0.273.
To test the accuracy of the empirical percentile $r_{0.1}$, we drew 216 samples without replacement from the data pool and repeated this procedure 5000 times. The choice of 216 samples will be justified later in Section 6.2. For each run, we computed the empirical percentile, then obtained the mean and standard deviation across all runs. From these, we calculated the coefficient of variation, which measures variation relative to the mean. The relative variation lies between $1.4\%$ and $10.9\%$, with the maximum observed for upward flexibility at hour 10 and the minimum for energy flexibility at hour 19, which is also the case where the Weibull distribution provides the best fit.
5 Analytical solution
This section presents the proposed analytical solution for addressing the chance-constrained bidding problem introduced in Section 3.
5.1 Bonferroni correction
The joint chance constraint in (2) can be decomposed into three individual constraints using the Bonferroni correction (BonferroniCorr). This adjustment redistributes the overall allowance for violation equally across each constraint, effectively lowering the significance level for each test to ensure that the total probability of violation remains within the desired limit:
@@ 5.1 Bonferroni correction
The Bonferroni approach essentially replaces the solution space for each decision variable in (2b) with a smaller set, as presented in (7), which may make the solution more conservative. On the other hand, one might replace (2b) with a larger solution space by setting the right-hand side of the inequalities in (7) to $\varepsilon$. These sets may act as lower and upper bounds on the solution of (2), respectively. We elaborate with a two-dimensional example: using the joint probability $\mathbb{P}(x\leq X,y\leq Y)\geq 1-\varepsilon$, shaded in gray in Figure 5, we can restate it as:
@@
Any solution of $x$ and $y$ which satisfies (8d) will then satisfy (8c) $\Rightarrow$ (8b) $\Rightarrow$ (8a). The inequality (8c) implies that threshold values of failure probability are distributed over any combination of $x$ and $y$, meaning that $x$ and $y$ may take any values satisfying $\mathbb{P}(X\leq x)<\epsilon$ or $\mathbb{P}(Y\leq y)<\epsilon$ as long as the sum of the two probabilities also satisfies (8c). The Bonferroni corrected version (8d) expresses that at most $\varepsilon/2$ of the failure can be associated with each variable, meaning that the possible values of $x$ and $y$ have a tighter constraint. If the distribution of $X$ and $Y$ are the same, while the intersect is zero, then this is the exact solution where (8d) = (8c). In Figure 5, this is the case if we assume that the probability mass in the pink and purple shaded areas (excluding the joint probability area) are equal, i.e., $\mathbb{P}(X\leq x)=\mathbb{P}(Y\leq y)$ and $\mathbb{P}\left(X<x,Y<y\right)=0$.
xxyy $\mathbb{P}(X>x,Y>y)$ $\mathbb{P}(Y\leq y)$ $\mathbb{P}(X\leq x)$ $\mathbb{P}(X\leq x,Y\leq y)$
5.2 Analytical reformulation of (2)
Following the procedure in Section 4, the Weibull distribution is fitted to data under a threshold level, $r_{0.1}$, which in the present case is the empirical 10th percentile of each flexibility, i.e., $\mathbb{P}(R\leq r_{0.1})=\varepsilon$. Applying Bayes’ theorem, we can thus express the conditional probability for an arbitrary random variable $R$ being under a certain threshold $b$ as:
@@ 5.2 Analytical reformulation of (2)
6 Sample-based solution for benchmarking
@@
6.1 Sample-based reformulation of the optimization problem
@@
where $\{M^{\uparrow},M^{\downarrow},M^{\rm{E_{20}}}\}$ are sufficiently large positive constants, and $y_{i}$ is an auxiliary binary variable corresponding to sample $i$. Constraints (13b)-(13d) ensure that if $y_{i}=0$, all constraints are fulfilled for a specific sample $i$, while if $y_{i}=1$, these constraints are relaxed. Constraint (13e) ensures that the total budget for constraint violations is preserved, allowing for at most a 10% violation of the samples, in accordance with the P90 requirement. Due to the reduced number of data points in the case study, the problem remains tractable and can be solved without the need for additional simplification or relaxation.
6.2 Sampling
To determine the minimum number of randomly selected samples necessary to accurately represent the estimated data, the following formula is used (SampleSizeSAA):
@@ 6.2 Sampling
6.3 Ex-post out-of-sample validation of constraints
@@
As a summary, the sampling, distribution fitting, decision making, and out-of-sample validation procedure is illustrated in Figure 7. The randomly sampled data points are used in both the analytical and sample-based approaches, with the former utilizing only a selected subset of the data points, as detailed in Section 4.
[…]
Figure 7: Flowchart of the sampling, distribution fitting, decision making, and out-of-sample (OOS) validation procedure. From the 366 available data points, 216 are randomly sampled without replacement. Among these, the 22 samples below the empirical 10th percentile are used to fit a Weibull distribution via maximum likelihood estimation. The estimated parameters, $\hat{\kappa}$ and $\hat{\gamma}$, are then assessed with a KS test against the same 22 samples. The remaining 150 samples are employed for out-of-sample analysis of bid violations under both the sample-based and analytical optimization approaches.
7 Numerical results
Our case study relies on real-world data comprising charging profiles from 1400 residential EV charging boxes in Denmark. The measurements were collected between March 24, 2022, and March 21, 2023, with an average time step of 2.84 minutes. To ensure a uniform resolution, the data were interpolated to one-minute granularity. It is assumed that each charging box corresponds to a single EV. The historical EV consumption levels serve as the baseline for estimating flexibility. Further details of the case study are provided in Appendix A.
In this section, we first present numerical results regarding the bidding strategy of the EV aggregator, obtained from both the analytical and sample-based approaches. We also report the rate of out-of-sample constraint violations and provide a sensitivity analysis with respect to $\varepsilon$. All source codes are publicly available in Git, while the historical EV consumption data remain proprietary. The estimated values of $\hat{\kappa}$ and $\hat{\gamma}$ from the distribution fittings are also reported in Git.
All optimization was performed on an HP EliteBook 840 14-inch G10 Notebook PC, equipped with a 13th Gen Intel® Core™ i7-1365U processor (10 cores, 12 logical processors), using Visual Studio Code. The sample-based optimization completes in 44.09 seconds, averaging 1.84 seconds per hour (10 runs with different sample splits). In comparison, the optimization based on the analytical formulation of the joint chance constraint runs in 9.19 seconds, averaging 0.38 seconds per hour over the same 10 runs, corresponding to a speed-up factor of 4.8 relative to the sample-based approach. When including the time required for distribution fitting, the full analytical procedure completes in 32.9 seconds, averaging 1.33 seconds per hour, which corresponds to a speed-up factor of 1.34–1.38 compared to the sample-based method.
Figure 8: The two top plots illustrate the capacity bids for FCR-D up (left) and FCR-D down (right) services obtained with the proposed analytical method (blue) and the benchmark sample-based method (pink). The remaining four plots show the out-of-sample constraint violation rates, reported separately for each constraint as well as in aggregate. All results represent the mean over 10 runs, where each run uses different in-sample and out-of-sample sets, while the number of samples in each set remains fixed.
7.1 Bidding strategy
As described earlier, we consider 10 runs, each based on 216 randomly selected in-sample data points drawn from the 366 available points. The top plots of Figure 8 illustrate the average reserve capacity bids across all runs for the FCR-D up and down markets. The remaining plots report the average constraint violation rates, obtained through the out-of-sample validation procedure outlined in Section 6.3, using 150 samples not included in the optimization.
As expected from the validation discussion in Section 4.3, the variation of the empirical percentile does not affect the upward bids, since no bids are placed in this direction for either hour. In the downward direction, however, the relative variation of the bids is 7.23% for hour 19 and 12.84% for hour 10. The larger relative variation of the bids at hour 19 may not correspond directly to the lower relative variation of the empirical percentile of energy flexibility at this hour, but could instead reflect the combined relative variation of the empirical percentiles across all flexibilities, which range between 1.4% and 4.4%.
@@ 7.1 Bidding strategy
where $\pi_{h}^{(\cdot)}$ represents the prices of the FCR-D capacity reservations, and $b^{(\cdot)*}_{h}$ denotes the bid placed at hour $h$ in either direction. Using prices from the early auction of the FCR-D up and down markets for the relevant period, corresponding to the data collection timeframe, and the resulting reserve bids, the total reservation payment amounts to €136,292 for the sample-based method and €105,366 for the analytical method over the entire year. Extending the formula to include penalties for bid violations would provide a more complete picture of the aggregator’s profit; however, this is outside the scope of the present paper. Nevertheless, considering the results regarding constraint violations, it is likely that bids determined using the sample-based method may incur higher penalties.
7.2 Sensitivity analysis
For the sensitivity analysis, we focus exclusively on the proposed analytical approach, investigating how the optimal reserve capacity bids vary with different values of $\varepsilon$. In particular, we examine the effect of smaller $\varepsilon$, which imposes stricter reliability requirements on the bids.
In (10), the allowance for failure is distributed evenly across all requirements involved in the chance constraints. We can utilize this formulation to create a constraint that allows for a failure allowance lower than the level used in the distribution fitting scheme. Recall that we are using data from the 10th percentile to fit distributions to each flexibility. This allows us to investigate whether we can find solutions to the bidding strategy with a violation allowance lower than 10%. To perform the sensitivity analysis on the failure allowance, we can define the constraints similarly to (12b)-(12d), replacing the right-hand side of (10) with an arbitrary $\alpha\in(0.0005,0.1]$, where $\alpha=\varepsilon/3$ corresponds to the Bonferroni corrected problem (12). Following the same procedure as in Section 5, we obtain a formulation of the optimization problem for the bidding strategy where we may vary the input reliability level, below the threshold $\varepsilon$: